• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name W. Deconinck

    Log entry time 23:07:11 on October 23, 2009

    Entry number 296379

    keyword=Status of the experiment (overview)

    During the swing shift, the following spectrometer related problems occurred:

    - The left arm Q2 power supply is unable to deliver a current of more
    than 500 A. This conclusion was established after 5 failed attempts to
    ramp up with trips at identical currents, and without effect from or on
    cryo parameters. The tech (Jack) who will be able to diagnose and fix
    this problem is out of town until Monday. There do not seem to be any
    problems with the other magnets in the left arm.

    - The right arm magnets could not be ramped up due to quenches or trips
    (from the effect on cryo they were likely real quenches). The reason was
    traced back to problems with the He level meters that show incorrect
    values. With Ed Folts we power-cycled these modules, and swapped some of
    them out, but the problems persisted even in the new modules. In effect,
    the false readings on the level meters cause the helium level to rise to
    100% and might be the cause of the quenches. Ed will come in again
    tomorrow to diagnose the problems more.

    Due to these problems we will be unable to go to (one-arm) production
    running until at least tomorrow day after Ed will have had a chance to
    check out more things on the right arm.

    We will be taking Compton runs overnight with an injector expert
    available at MCC. There are however some vacuum problems in the BSY Hall
    A beamline that are preventing us from starting that before 11pm.

    A copy of this log entry has been emailed to: paschke, mcnulty, rom, kees